OpenStudy (anonymous):

What is the 45th term of the arithmetic sequence 16, 20, 24, 28, ..

OpenStudy (anonymous):

Well, each new term is 4 more than the last one, so start adding. Or you can do this: \[a_n = a_1 + (n – 1)d\] Where a_1 is the first term in the sequence, d is the number they are increasing by, and n is the term number you want: \[a_{45} = 16 + (45 – 1)4 = 16+44*4 = 176+16=192\]

OpenStudy (anonymous):

So the 45th term is 192

OpenStudy (anonymous):

Thanks..

OpenStudy (anonymous):

No prob, mind marking best response?

OpenStudy (anonymous):

got cha!